OpenStudy (anonymous):

The three Math Idol judges have been eliminating contestants all day! The number of one-step equations and two-step equations who have been eliminated today is equal to 1070! If three times the number of one-step equations minus twice the number of two-step equations is equal to 110, how many one-step equations auditioned today? 900 620 520 450

OpenStudy (anonymous):

450 is what I got. x+y = 1070 3x-2y = 110 Multiply first equation 2x + 2y = 2140 Now add to second equation You get 5x = 2250 x = 450